OpenStudy (anonymous):

can someone helps??? please! Y varies inversely with x. If y = 6 when x = 5, find x when y = 7

OpenStudy (whpalmer4):

if y varies inversely with x, that means that \(y = \frac{k}{x}\) where \(k\) is a constant. Use y = 6 and x = 5 to find \(k\), then use \(k\) to find the value of x when y = 7.

OpenStudy (johnweldon1993):

Right, as whpalmer has shown \[y = \frac{ k }{ x }\] putting your values to solve for k \[6 = \frac{ k }{ 5 }\] how do we solve for 'k'? we multiply both sides by 5 k = 30 now you go back to the original equation \[y = \frac{ k }{ x }\] you have a 'y' value and now a 'k' value...so plug those in and solve for 'x' \[7 = \frac{ 30 }{ x }\] x = ...?